Integrate the following kindly please give me answer.. x^2 e^x^3 cos(e^x^3)dx

Grade:12

1 Answers

Arun
6 years ago
Assume 
(e^x³) = t
Then
e^x³ * 3x² dx = dt
Hence
e^x³ * x²dx = dt/3
Question becomes
(1/3) Cos t dt
(1/3) sint + C
Now put the value of t
(1/3) sin( e^x³) + C